A projectile leaves the ground at an angle 30° to the horizontal. The kinetic energy is E neglecting the air resistance . Find in terms of E it's kinetic energy at the highest point of motion




1
Expert's answer
2022-02-01T09:50:47-0500

At highest point only vertical components of velocity

"v_x=vcos\theta=vcos30°=\frac{\sqrt3}{2}"

"E_{up}=\frac{1}{2}mv_x^2\\E_{up}=\frac{1}{2}mv^2cos^2\theta"

"E_{up}=\frac{1}{2}mv^2\times(\frac{\sqrt{3}}{2})^2=\frac{3}{8}mv^2"

"E=\frac{1}{2}mv^2"

"\frac{E_{up}}{E}=\frac{\frac{3}{8}mv^2}{\frac{mv^2}{2}}"

"E_{up}=\frac{3}{4}E"